vidyo

New vidyo room:

(old vidyo room: vidyo link PIN 1055)

News:

Roberto: contacted by M. Cruz (muon manager candidate), saying that in case that she is elected she'll move to 100% on the position (no opinion on her). The other candidate is ok. We need to decide who to vote. For the DT there is only one candidate (Ignacio), nobody from IT. Ignacio is very competent but quite firm on his opinions. The new managers will have a lot of influence on the decisions for the newt detector.

Tommaso: seminare on june 6 afternoon on ML for uncertanties evaluation.

Framework & production

Federica: access for latino dataset only via xroot

Analysis

B physics

Martino: tomorrow on London for parking brainstorming. Starting on the preparations of Run3 (alessio involved in trigger, paolo in studies for high PU). Meeting on 28th May, few presentations planned (possible LFV studies, possible use of PU as B source).

Alessio: still on efficiencies for P'5

Alberto: starting big production (final or next to final ones)

Heavy New Physics

All mc signal samples submitted, working on systematics for 2017/2018.

VBS

News on the Latino 2016 miniAODs transfer from Eos to Legnaro: first sample successfully copied with webfts. Waiting for feedback from Latino group before proceeding.

VH SM

1.) Background Study:

a.) 2016 dataset + MC complete (private NANOAODv0)

i.) Running in analysis production

b.) 2017 dataset + MC (Nanoaodv0)

i.) Missing MC Phdex transfer complete. (https://cmsweb.cern.ch/phedex/prod/Data::Subscriptions#state=create_since%3D1555749958%3Bnode%3D45%3Brequest%3D1755644), but missing two samples

c.) 2017 dataset + MC (Nanoaodv4)

i.) in progress

2.) Signal Study:

a.) private produced LO WHWW (NANOAOD) complete (2016 setting)

i.) signal acceptance/sensitivity (trigger,event selection) study : Will be sending out to Ugo and Marco today.

b.) private produced NLO multijet matched WHWW (2016 setting)

i.) Fail at first production chain, troubleshooting

3.) Analysis code:

a.) reviews on the cross-cleaning , mother-daughter ID module, ...

b.) implement lep SF, trigger SF (ISOmu24 OR ISOTRKMu24), no ele trigger SF, GEN weight, PU weight (benchmark at 2016)

4.) TnP study on ele trigger (HOLD)

Same sign dileptons

EPR

Siewyan (Sherpa generator study)

Mia: possible EPR under traker

1.) Heavy Flavour

a.) produce corresponding Sherpa sample for ZBJetsToNuNu and ZJetsToNuNu _BGenFilter (FALL17) using b enrichment filter in SHERPA before moving to LO.

2.) Exotic Physics (Dark photon)

a.) Is SHERPA capable of auto width compute as madgraph? in progress

Trigger(less)

Remove parenthesis from next meeting.

Detector

DT

Timing Layer

Paolo: friday afternoon discussion between italians Martino: activity starting on possibilities for bph, M. Dinardo (milano Bicocca) working maybe with a students (maybe a Master thesis), probably look into P'5.

Tracker

Nothing to report.

A.O.B.

-- SiewYanHoh - 2019-05-20

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2019-05-20 - AlbertoBragagnolo
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback